Estimate the maximum core diameter for an optical fiber with refractive index difference of 1.45% and core refractive index of 1.52 in order that it may be suitable for single-mode operation when operating wavelength is 0.85μm. Also calculate cut off wavelength cl if core diameter is 1.1μm. 


1
Expert's answer
2021-11-24T19:13:57-0500

"d=\frac{V\lambda}{\pi n\sqrt{2\Delta}}=0.43~\mu m,"

"\lambda_1=\frac{\pi d_1n\sqrt{2\Delta}}{V}=2.2~\mu m ."